# Word length predicts word order: “Min-max”-ing drives language evolution

### v1

@@ 2.2.3 Relationship between word order and noun/verb lengths
To see whether word order is related to the length of nouns and verbs I conducted a repeated-measures ANOVA with the average lengths of nouns and verbs as the repeated measures and the word order (SV, VS, free) as the between subject factors. Analysis of the results of word order and its relation to length of the different argument and predicate categories is quite striking. First, both SV and VS languages have a highly significant effect of word length, [^10] Additional statistical comparisons and plots are available at the linked repository (Fn 4). whereby arguments are, overall, shorter than predicates crosslinguistically (Figure 1; p < 0.001). [^11] It is also worth noting here that in this corpus words are overall shorter in SV languages than VS languages, which in turn are shorter than those in ‘free’ languages. This displays a general observed trend that languages with free word order tend to have more morphology to identify argument roles and other grammatical relations (Koplenig et al. 2017; Yadav et al. 2020).
@@
If we consider only nouns and only verbs, the findings are similar when we normalize the dataset by reducing the vocabulary to unique items. However, if we allow for frequency effects (i.e. we do not reduce the vocabulary to unique items) there is a clear correlation with word order: nouns are longer than verbs in SV languages, but shorter than verbs in VS languages (Figure 2; p < 0.001). This is a major finding that we will return to in the discussion. Setting it aside for now, we turn to two different tests of this measure relating to diachrony: predicting word order change, and the potential interaction with family membership (descent) and area (geography).
2.3 Testing the predictive validity of word length
@@
Given the observation that relative lengths of nouns and verbs differ in corpora for languages with different word orders, we can ask how predictive this observation is. That is, can we predict the word order of a language based on the relative lengths of its nouns and verbs? If we are strictly interested in identifying basic word order from corpora, it is clear that the N1 ratio already allows us to do so. However, we are also interested in understanding how noun and verb lengths may be related to word order. Being able to predict word order from noun and verb lengths in a corpus for both a modern language and its ancestral language would indicate that such properties are related to how word order evolves across time.
@@
To investigate this question, I sourced additional corpora of tagged data from two pairs of historical and modern languages in which the modern language changed its word order: 1) Ancient Hebrew (VS) and Modern Hebrew (SV) and 2) Classical Arabic (VS) and Egyptian Arabic (SV). All languages belong to the Semitic branch of the Afroasiatic language family, a necessary constraint given the limited amount of data available for most pre-modern languages. In the case of Hebrew, the UDT has POS-tagged sentences for both Ancient [^12] https://universaldependencies.org/treebanks/hbo_ptnk and Modern [^13] https://universaldependencies.org/treebanks/he_htb varieties. In the case of Classical Arabic I used the Quranic Arabic Corpus [^14] Version 0.4: https://corpus.quran.com and for Egyptian Arabic I used the BOLT Egyptian Arabic Treebank (Maamouri et al. 2018).151515(Discussion Forum) https://catalog.ldc.upenn.edu/LDC2018T23 If the lengths of nouns and length of verbs in these corpora allows us to correctly classify Ancient Hebrew and Classical Arabic as VS and the modern varieties as SV, this would suggest that these are salient properties affecting how word order has developed diachronically. Importantly, given the textual differences between the tagged PBC (New Testament texts) and the test corpora (Biblical, Quranic, modern, and informal texts), such a finding would be significant.
@@
In order to conduct this test, the data for Ancient Hebrew (ISO 639-3: hbo), Modern Hebrew (ISO 639-3: heb), Classical Arabic (cla) [^16] Classical Arabic does not have an assigned ISO 639-3 code or Glottocode. Accordingly, I use the abbreviation cla here, as there is no current language assigned to this code in the ISO 639-3 tables. and Egyptian Arabic (ISO 639-3: arz) were processed and the respective word classes were counted and measured as per the method outlined for the tagged PBC (each corpus > 400 sentences, see Table 1). The average length of nouns and verbs (including frequency effects) were then used as continuous features for a Gaussian Naive Bayes (GNB) classifier.
@@
Using these five features, a GNB classifier was trained on the original combined word order dataset containing hand-coded classification of 974 languages (Modern Hebrew and Modern Arabic were removed), and then used to predict word order for our pairs of languages: Ancient Hebrew and Modern Hebrew on the one hand, Classical Arabic and Egyptian Arabic on the other. Ancient Hebrew was classified as VS while Modern Hebrew was classified as SV by the trained classifier. Classical Arabic was classified as VS while Egyptian Arabic was classified as SV. This indicates that the relative lengths of nouns and verbs, when frequency is taken into account, can predict word order from corpora, with various implications for our understanding of language processing and language change. This is a striking finding, and I am unaware of any other research that has been able to show predictive validity of a model related to word order change. However, it is possible that this result is conditioned to some degree by the fact that these are Semitic languages spoken in the same geographical area, so we now turn to an investigation of whether descent or location provides a better explanation for the correlation of noun and verb lengths with word order.
2.4 Testing variance predicted by word length
Dunn et al. (2011) showed that descent from a common ancestor was a better predictor of many grammatical patterns than other explanations. Although Jäger & Wahle (2021) counter this finding with additional data, we can also assess this claim using the current dataset to see if descent is more explanatory of word order for a given language than the lengths of nouns and verbs. Additionally, in more recent work, Naranjo & Becker (2022) show that areal affects are as important as family membership in predicting linguistic phenomena, finding that the results in Dunn et al can be just as easily predicted by language area. This is also a claim that we can test with our dataset. Accordingly, this investigation can proceed with several (motivated) simplifying observations.
@@ 2.4 Testing variance predicted by word length
In particular, a binary dependent variable for word order means we can use hierarchical linear regression, which allows us to observe the relative impact of particular features that are theoretically relevant to the question at hand by seeing how they influence the performance of successive models. In the case of word order, we can train a regression model with a feature known to be relevant to word order (the N1 ratio), and then subsequent models that incorporate additional features of interest (“language area/location”, “family membership”, “noun/verb lengths”). We can then compare the models to see whether the additional features increase the variance explained in the data via the F-value: an increase in F-value (change) and the significance of this change indicates more/additional variance explained by the particular feature added in a model.
Accordingly, for all languages in the tagged PBC I retrieved the family membership and geographical information from Glottolog, filling in missing data via Ethnologue and other sources. Then I filtered the data on four conditions: a) languages found in Dunn et al that are also present in the tagged PBC (106 languages), b) languages found in the tagged PBC that are in the four families investigated by Dunn et al (539 languages), c) languages in families with more than 75 members in the tagged PBC (962 languages), and d) languages in families with 2 or more members in the tagged PBC (1,486 languages). For each of these conditions, I ran a hierarchical linear regression, adding features of interest progressively to see how they impacted the fit of the successive models and whether particular features account for additional variance.
The base model (Model 1) uses the N1 ratio as the primary feature for classification. This is because we know from comparison with expert determinations of word order (see footnote 5 above) that the N1 ratio significantly correlates with the classification of languages as SV/VS. Due to the strong relationship observed between language area and typological properties of language (see Naranjo & Becker 2022), Model 2 then includes GPS coordinates and macroarea identifiers for each language as features. Model 3 additionally incorporates descent (language family membership) as a feature (as per Dunn et al. 2011), and Model 4 includes the average lengths of nouns and verbs (as per the relationship identified in this paper). If descent is more explanatory than the lengths of nouns/verbs, this should be clear from the differences between the respective models for each condition. Results are displayed in Table 2.
@@
1 (N1 ratio) 106 45.85 7.8e-10
2 (Lg area) 106 20.40 2.4e-12 8.58 3.9e-05**
3 (Family) 106 16.70 5.8e-12 1.48 0.227
4 (N/V len) 106 13.40 4.9e-12 3.27 0.042*
@@
1 (N1 ratio) 539 230.41 1.4e-43
2 (Lg area) 539 100.00 1.7e-63 39.86 2.9e-23**
3 (Family) 539 80.71 5.3e-63 2.46 0.118
4 (N/V len) 539 64.75 3.5e-67 14.57 6.9e-07**
@@
1 (N1 ratio) 962 426.74 9.9e-79
2 (Lg area) 962 220.49 4.2e-134 105.35 6.1e-59**
3 (Family) 962 176.22 6.6e-133 0.04 0.839
4 (N/V len) 962 130.93 8.3e-135 9.69 6.7e-05**
@@
1 (N1 ratio) 1486 586.46 1.8e-109
2 (Lg area) 1486 193.61 1.3e-133 45.19 6.0e-28**
3 (Family) 1486 155.10 1.2e-132 1.03 0.309
4 (N/V len) 1486 115.43 3.5e-135 11.02 1.7e-05**
@@
From these analyses we can see that, as expected, the N1 ratio significantly predicts word order under all conditions. Additionally, language area is a significant predictor of word order that adds to the strength of the classification. Language family (descent) does not add additional predictive power to our model under any conditions (over language area), whereas the length of nouns and verbs does. This analysis shows that the length of nouns and verbs in a language accounts for more variance and at the same time is more significant than that language’s family membership in determining its word order classification, under all test conditions. When we remove language area from the linear regression this is still the case, [^18] See additional statistical results at the linked repository (Fn 4). indicating that while language area is a stronger predictor than family membership, lengths of nouns and verbs is yet a stronger one.
This largely aligns with Jäger and Wahl’s (2021) finding regarding the (lower) correlation between descent and linguistic structures in comparison to other explanations, with the caveat that Jäger and Wahl included language isolates, which is likely to greatly diminish the affect of descent. If a large number of languages in a sample are not related to others, clearly descent can have only a minimal impact on word order. By including only language families with two or more members (which excludes 111 languages in the taggedPBC), I show that descent does not explain much variance above and beyond language location/area, in contrast with noun and verb lengths. This also suggests that language family/descent is highly correlated with linguistic area, which conforms with expectations related to human history and migration. Additionally, the amount of variance explained seems to depend on the number of languages (and language families) in the sample, highlighting the need for more data for languages from under-represented language families.
@@ 3 Discussion
That arguments overall are shorter in languages of the tagged PBC, across all word orders, suggests that there are certain (non-Noun) arguments affecting our observations. Particularly, the effect of PRON seems quite likely, whereby a tendency for pronouns to be shorter crosslinguistically, along with a preference for pronouns as arguments, may impact the observation that arguments are shorter overall than predicates. This pattern aligns with the processing claim that ‘heavier’ words (or word classes in this case) occur later in a sentence, partly validating Hypothesis 1. The counter-argument, also from the data, is that nouns are longer than verbs in SV languages and shorter than verbs in VS languages when frequency is considered, validating Hypothesis 2. This suggests that speakers prefer to place heavier (more “contentful”, or perhaps “novel”) constituents before lighter ones crosslinguistically in many cases.
These seemingly antithetical results partially validate both the “weight”-based theory and the “surprisal” theory, highlighting the interaction of two pressures: one whereby speakers prioritize lighter constituents (possibly for the purposes of production/processing efficiency), and one where they prioritize heavier constituents (possibly for informational purposes). The two perspectives may in fact be looking at the same data from different angles. On the one hand, grouping all arguments into a single class highlights the overall impact of shorter strings, since the most frequently used arguments (pronouns, demonstratives) are typically shorter crosslinguistically. On the other, more “content-ful” word types that aid planning and inference or predictability (nouns, verbs) will typically be longer crosslinguistically.
@@ 3.1 A “Min-Max” theory of processing
The current proposal re-formulates or refines Hawkins’ PGCH to incorporate (increased) complexity as a motivation for the order of constituents, bearing some resemblance to proposals in Levshina (2023). I frame this as (to borrow a term from computer gaming) the “Min-Max” theory of processing (schematized in Figure 3). The idea here is that speakers seek to minimize their processing effort while maximizing the information in their utterances. These are competing/parallel pressures that speakers are constantly balancing in production (or “performance”) in order to choose the best strategy for their immediate goals. As speakers produce variants along a cline of grammaticality, these variants compete on the basis of multiple factors, including how well such variants support Min-Max-ing in relation to other variants, leading to language change.
@@
The existence of multiple competing pressures affecting the development of language, as described by this Min-Max theory, is supported by work in clinical psychology (e.g. Rezaii et al. 2023) and language modeling (e.g. Hahn & Xu 2022), as well as findings from information theory (Zaslavsky 2020; Tucker et al. 2025). In a study of patients with nonfluent aphasia, Rezaii et al. (2023) showed that patients with this condition compensate for their deficit in producing complex grammatical structures by increasing the information content of their utterances, dropping high-frequency (function) words in favor of low-frequency (content) words. By modeling variations in grammatical structure based on corpora from 80 languages, Hahn & Xu (2022) found that the resulting models were less optimized for dependency length and information locality than the original structures of each language, indicating that languages evolve to efficiently balance the two considerations. [^20] Notably, Hahn & Xu propose a theory that the interaction between Dependency Length (DL) and Information Locality (IL) accounts for word order variation and change, but neither of the two concepts, as expressed, make claims about directionality/order, only about the relationships between (internal) constituents. In contrast, the Min-Max theory makes the claim that speakers balance information content (as reflected by longer nouns in SV languages and longer verbs in VS languages) with processing/production concerns, and that this interaction is what drives the evolution of word order. As nouns lengthen or verbs shorten in a language, this will affect their placement in a stream, which then impacts the overall order observed in surface structures. That is, the Min-Max theory makes a universal claim regarding language-external pressures, while DL and IL are language-internal concepts. These and other studies indicate that language users are sensitive to efficiency as it pertains both to processing and information content.
@@
The sensitivity of language users to processing efficiency and information content has also been a major topic of investigation in information theory, which is concerned in part with the encoding and decoding of information streams. One information-theoretic proposal is that semantic systems evolve following an “Information Bottleneck” principle (IB; Tishby et al. 1999), whereby “optimal representations satisfy a tradeoff between compressing X, i.e. minimizing the representational complexity, and maximizing the relevant information about Y” (Zaslavsky 2020: 14). Work by Zaslavsky shows that IB can account for the development of color naming systems in language among other kinds of sytems, and additional research shows the utility of IB in machine learning and distributional models of semantics (Slonim & Tishby 2000).
@@
The implications of this principle have most recently been examined in agent-based emergent learning simulations with the addition of a “utility” principle (Tucker et al. 2025) – the interaction between “utility”, “informativeness”, and “complexity” allows more human-like communication to emerge among agent models in their study. Such a finding is very much supported by the results reported here, though I would argue that the Min-Max theory offers a more parsimonious account of human behavior as related to language specifically. While the current corpus evidence applies at the word level in relation to basic word order, it seems likely that the minimization of processing difficulty would interact with the maximization of information content at multiple (hierarchical) levels. Speakers min-max in order to select words in a stream (“informativeness” and “complexity” as per Tucker et al.), while simultaneously min-maxing to select relevant structures related to their communicative goals (Tucker et al.’s “utility”).
@@
The Min-Max theory also bears some resemblance to Levshina’s (2023) proposals, described in terms of communicative efficiency and accessibility. However, such indicators cannot fully explain certain patterns in grammar such as “harmonic” word order. A recent proposed explanation for such patterns involves frequency of word classes, whereby Mansfield & Krapp (2025) suggest that “harmonic ordering is really a frequency effect: as replication converges on a phrase structure with a consistent linear order, the most frequent word class tends to be at one edge.” Although they did not test token length in their model, it is notable that in the current study frequency is a major factor in identifying the relationship between length and word order, suggesting that the most frequent word class in their model may also be the longest element of the domain they are testing, which would be driven by a maximization of information principle. The Min-Max theory also seems to provide an explanation for non-parallel (hierarchical) harmonic word order, which is not explained by their model.
@@
The Min-Max theory may additionally provide an explanation for why word forms remain long (or lengthen), as described for particular languages (see Cristofaro 2021; Haspelmath 2021; Petré 2017). Here, again, the maximization of information principle would support the development or retention of particular (long) word forms, as speakers balance this need against the minimization of processing difficulty. As noted above, this proposal is supported by evidence from color-naming systems, which show patterns in keeping with the Information Bottleneck principle. The Min-Max theory would suggest that this principle does not only hold for “forms” (of, i.e. color words), but also governs “behavior”: speakers behave in ways that are consistent with maximizing information in streams (via word order), while minimizing processing difficulty.
@@
As noted in section 2, such a theory can be conceived of as a universal (underlying) pressure that applies generally to how language is realized. From an innate perspective, speaker choices impacted by such pressures would lead to a re-analysis or re-formation of rules that guide surface realizations. From a usage-based perspective such pressures would lead to the development of new surface variants. These are not necessarily opposing views (see Rastelli 2025), and the outcomes are quite similarly dependent on language being viewed as a complex adaptive system, whereby diachronic changes may have a combination of both “result-oriented” and “source-oriented” explanations (see Cristofaro 2019; Schmidtke-Bode & Grossman 2019).

### v2

@@ 2.1.3 Relationship between word order and noun/verb lengths
To see whether word order is related to the length of nouns and verbs I conducted a repeated-measures ANOVA with the average lengths of nouns and verbs as the repeated measures and the word order (SV, VS, free) as the between subject factors. Analysis of the results of word order and its relation to length of the different argument and predicate categories is quite striking. First, both SV and VS languages have a highly significant effect of word length, [^9] Additional statistical comparisons and plots are available at the linked repository (Fn 2). whereby arguments are, overall, shorter than predicates crosslinguistically (Figure 3; p < 0.001). [^10] It is also worth noting here that in this corpus words are overall shorter in SV languages than VS languages, which in turn are shorter than those in ‘free’ languages. This displays a general observed trend that languages with free word order tend to have more morphology to identify argument roles and other grammatical relations (Koplenig et al. 2017; Yadav et al. 2020).
@@
If we consider only nouns and only verbs, the findings are similar when we normalize the dataset by reducing the vocabulary to unique items. However, if we allow for frequency effects (i.e. we do not reduce the vocabulary to unique items) there is a clear correlation with word order: nouns are longer than verbs in SV languages, but shorter than verbs in VS languages (Figure 4; p < 0.001). This is a major finding that we will return to in the discussion. Setting it aside for now, we turn to two different tests of this measure relating to diachrony: predicting word order change, and the potential interaction with family membership (genealogy/descent) and area (geography).
2.2 Study 2: Predictive validity of word length
@@
Given the observation that relative lengths of nouns and verbs differ in corpora for languages with different word orders, we can ask how predictive this observation is. That is, can we predict the word order of a language based on the relative lengths of its nouns and verbs? If we are strictly interested in identifying basic word order from corpora, it is clear that the N1 ratio already allows us to do so. However, we are also interested in understanding how noun and verb lengths may be related to word order. Being able to predict word order only from noun and verb lengths in a corpus for both a modern language and its ancestral language would indicate that such properties are related to how word order evolves over time.
To investigate this question, for Study 2 I sourced additional corpora of tagged data from two pairs of historical and modern languages in which the modern language changed its word order:
  1. 1.
    Ancient Hebrew (VS) and Modern Hebrew (SV)
  2. 2.
    Classical Arabic (VS) and Egyptian Arabic (SV)
@@
All languages belong to the Semitic branch of the Afroasiatic language family, a necessary constraint given the limited amount of data available for most pre-modern languages. In the case of Hebrew, the UDT has POS-tagged sentences for both Ancient [^11] https://universaldependencies.org/treebanks/hbo_ptnk and Modern [^12] https://universaldependencies.org/treebanks/he_htb varieties. In the case of Classical Arabic I used the Quranic Arabic Corpus [^13] Version 0.4: https://corpus.quran.com and for Egyptian Arabic I used the BOLT Egyptian Arabic Treebank (Maamouri et al. 2018). [^14] https://catalog.ldc.upenn.edu/LDC2018T23 (Discussion Forum) If the lengths of nouns and verbs in these corpora allows us to correctly classify Ancient Hebrew and Classical Arabic as VS and the modern varieties as SV, this would suggest that these are salient properties affecting how word order has developed diachronically. Importantly, given the textual differences between the taggedPBC (New Testament texts) and the test corpora (Biblical, Quranic, modern, and informal texts), such a finding would be significant.
@@
In order to conduct this test, the data for Ancient Hebrew (ISO 639-3: hbo), Modern Hebrew (ISO 639-3: heb), Classical Arabic (cla) [^15] Classical Arabic does not have an assigned ISO 639-3 code or Glottocode. Accordingly, I use the abbreviation cla here, as there is no current language assigned to this code in the ISO 639-3 tables. and Egyptian Arabic (ISO 639-3: arz) were processed and the respective word classes were counted and measured as per the method outlined for Study 1 (each corpus > 400 sentences, see Table 1). The average length of nouns and verbs (including frequency effects) were then used as continuous features for a Gaussian Naive Bayes (GNB) classifier.
@@
Using these five features (nounlen, verblen, nlonger, nlen-vlen, nlen/vlen), a GNB classifier was trained on the original combined word order dataset from the taggedPBC containing hand-coded classification of 1,150 languages (Modern Hebrew and Modern Arabic were removed), and then used to predict word order for our pairs of languages: Ancient Hebrew and Modern Hebrew on the one hand, Classical Arabic and Egyptian Arabic on the other. Ancient Hebrew was classified as VS while Modern Hebrew was classified as SV by the trained classifier. Classical Arabic was classified as VS while Egyptian Arabic was classified as SV.
As an additional check, I sourced data from another historical language, Old Irish, as represented in the Codex Paulinus Wirziburgensis, whereby Latin text of St. Paul’s epistles was annotated by Irish scholars during the middle of the eighth century. The dataset (Table 2; Doyle 2018) contains over 3,600 Irish glosses/sentences. As an Indo-European language, Old Irish is widely surmised to have changed its word order to verb-initial from an earlier state, so this check allows us to verify whether the same properties hold in relation to lengths of nouns and verbs. Indeed, the classifier accurately classified Old Irish as VS, based on the data extracted from the Würzburg Glosses.
Language ISO 639-3 Data Source Sentences Unique nouns Unique verbs
Old Irish sga Würzburg 3,647 743 435
Table 2: Irish corpus used for testing historical word order change
Study 2 thus indicates that the relative lengths of nouns and verbs, when frequency is taken into account, can predict word order from corpora unrelated to the type of corpora contained in the taggedPBC, with various implications for our understanding of language processing and language change. This is a striking finding, and I am unaware of any other research that has been able to show predictive validity of a model related to word order change. However, it is possible that this result is conditioned to some degree by the fact that the majority of the test cases are Semitic languages spoken in the same geographical area, so we now turn to an investigation of whether genealogical descent or location provides a better explanation for the correlation of noun and verb lengths with word order.
2.3 Study 3: Testing variance predicted by word length
Dunn et al. (2011) showed that descent from a common ancestor was a better predictor of many grammatical patterns than other explanations. Although Jäger & Wahle (2021) counter this claim with additional data, we can also use the current dataset to see if descent is more explanatory of word order for a given language than the lengths of nouns and verbs. Additionally, in more recent work, Guzmán Naranjo & Becker (2022) show that areal affects are as important as family membership in predicting linguistic phenomena, finding that the results in Dunn et al can be just as easily predicted by language area. This is also a claim that we can test with our dataset. Accordingly, Study 3 can proceed with several (motivated) simplifying observations.
@@ 2.3 Study 3: Testing variance predicted by word length
In particular, a binary dependent variable for word order means we can use hierarchical linear regression (HLR), which allows us to observe the relative impact of particular features that are theoretically relevant to the question at hand by seeing how they influence the performance of successive models. In the case of word order, we can train a regression model with a feature known to be relevant to word order (the N1 ratio), and then subsequent models that incorporate additional features of interest (“language area/location”, “family membership”, “noun/verb lengths”). We can then compare the models to see whether the additional features increase the variance explained in the data via the F-value: an increase in F-value (change) and the significance of this change indicates more/additional variance explained by the particular feature added in a model.
Accordingly, for all languages in the taggedPBC I retrieved the family membership and geographical information from Glottolog, filling in missing data via Ethnologue and other sources. Then I filtered the data on four conditions: a) languages found in Dunn et al that are also present in the taggedPBC (120 languages), b) languages found in the taggedPBC that are in the four families investigated by Dunn et al (662 languages), c) languages in families with more than 75 members in the taggedPBC (1182 languages), and d) languages in families with 2 or more members in the taggedPBC (1,805 languages). For each of these conditions, I ran an HLR, adding features of interest progressively to see how they impacted the fit of the successive models and whether particular features account for additional variance.
The base model (Model 1) uses the N1 ratio as the primary feature for classification. This is because we know from comparison with expert determinations of word order (see Ring 2025) that the N1 ratio significantly correlates with the classification of languages as SV or VS. Due to the strong relationship observed between language area and typological properties of language (see Guzmán Naranjo & Becker 2022), Model 2 then includes GPS coordinates and macroarea identifiers for each language as features. Model 3 additionally incorporates descent (language family membership) as a feature (as per Dunn et al. 2011), and Model 4 includes the average lengths of nouns and verbs (as per the relationship identified in this paper). If descent is more explanatory than the lengths of nouns/verbs, this should be clear from the differences between the respective models for each condition. Results are displayed in Table 3.
@@
1 (N1 ratio) 120 45.27 6.4e-10
2 (Lg area) 120 24.29 1.3e-14 12.78 2.87e-07 **
3 (Family) 120 20.02 2.7e-14 2.05 0.154
4 (N/V len) 120 15.75 2.9e-14 3.17 0.045 *
@@
1 (N1 ratio) 662 189.99 3.6e-38
2 (Lg area) 662 95.65 4.1e-64 50.07 3.7e-29 **
3 (Family) 662 78.23 2.4e-64 5.79 0.016 *
4 (N/V len) 662 65.43 2.6e-71 21.30 1.0e-09 **
@@
1 (N1 ratio) 1182 378.83 2.1e-73
2 (Lg area) 1182 210.38 3.3e-136 116.99 2.6e-66 **
3 (Family) 1182 169.09 1.3e-135 2.72 0.099
4 (N/V len) 1182 127.47 2.3e-139 14.03 9.4e-07 **
@@
1 (N1 ratio) 1805 535.41 6.1e-104
2 (Lg area) 1805 178.89 3.8e-129 46.52 5.8e-29 **
3 (Family) 1805 143.49 2.3e-128 1.66 0.197
4 (N/V len) 1805 108.16 2.2e-132 14.46 5.8e-07 **
@@
From these analyses we can see that, as expected, the N1 ratio significantly predicts word order under all conditions. Additionally, language area is a significant predictor of word order that adds to the strength of the classification. Language family (descent) does not add additional predictive power to our model under most conditions (over language area), whereas the length of nouns and verbs does. This analysis shows that the length of nouns and verbs in a language accounts for more variance and at the same time is more significant than that language’s family membership in determining its word order classification, under all test conditions. When we remove language area from the linear regression this is still the case, [^17] See additional statistical results at the linked repositories (Fn 2; Fn 5). indicating that while language area is a stronger predictor than family membership, lengths of nouns and verbs is yet a stronger one.
This largely aligns with Jäger and Wahl’s (2021) finding regarding the (lower) correlation between descent and linguistic structures in comparison to other explanations, with the caveat that Jäger and Wahl included language isolates, which is likely to greatly diminish the affect of descent. If a large number of languages in a sample are not related to others, clearly descent can have only a minimal impact on word order. By including only language families with two or more members in the taggedPBC (which excludes 111 languages), I show that descent does not explain much variance above and beyond language location/area, in contrast with noun and verb lengths. This also suggests that language family/descent is highly correlated with linguistic area, which conforms with expectations related to human history and migration. Additionally, the amount of variance explained seems to depend on the number of languages (and language families) in the sample, highlighting the need for more data for languages from under-represented language families.
@@ 3 Discussion
That arguments overall are shorter in languages of the taggedPBC, across all word orders, suggests that there are certain (non-Noun) arguments affecting these observations. Particularly, the effect of PRON seems quite likely, whereby a tendency for pronouns to be shorter crosslinguistically, along with a preference for pronouns as arguments, may impact the observation that arguments are shorter overall than predicates. This pattern aligns with the processing claim that “heavier” words (or word classes in this case) occur later in a sentence, partly validating Hypothesis 1 of study 1. The counter-argument, also from the data, is that nouns are longer than verbs in SV languages and shorter than verbs in VS languages when frequency is considered, validating Hypothesis 2 of study 1. This suggests that speakers prefer to place heavier (more “contentful”, or perhaps “novel”) constituents before lighter ones crosslinguistically in many cases.
These seemingly antithetical results partially validate both the “efficiency”-based theory and the “surprisal” theory, highlighting the interaction of two pressures: one whereby speakers prioritize lighter constituents (possibly for the purposes of production/processing efficiency), and one where they prioritize heavier constituents (possibly for informational purposes). The two perspectives may in fact be looking at the same data from different angles. On the one hand, grouping all arguments into a single class highlights the overall impact of shorter strings, since the most frequently used arguments (pronouns, demonstratives) are typically shorter crosslinguistically. On the other, more “content-ful” word types that aid planning and inference or predictability (nouns, verbs) will typically be longer crosslinguistically.
@@ 3.1 A “Min-Max” theory of processing
The current proposal re-formulates or refines Hawkins’ PGCH to incorporate (increased) complexity as a motivation for the order of constituents, bearing some resemblance to proposals in Levshina (2023). I describe this as (to borrow a term from computer gaming) [^19] A “min-maxer” in gaming terminology is an individual who minimizes their effort while maximizing their gains. This can be in terms of focusing their efforts on activities in a game that give the best returns, or by minimizing survival stats on a character in order to maximize damage (as in a “glass cannon” build for ranged attackers in an rpg). the “Min-Max” theory of processing (schematized in Figure 5), whereby speakers seek to minimize their processing effort while maximizing the information in their utterances. These are competing/parallel pressures that speakers are constantly balancing in production (or “performance”) in order to choose the best strategy for their immediate goals. As speakers produce variants along a cline of grammaticality, these variants compete on the basis of multiple factors, including how well such variants support Min-Max-ing in relation to other variants, leading to language change.
@@
The existence of multiple competing pressures affecting the development of language, as described by this Min-Max theory, is supported by work in clinical psychology (e.g. Rezaii et al. 2023) and language modeling (e.g. Hahn & Xu 2022), as well as findings from information theory (Zaslavsky 2020; Tucker et al. 2025). In a study of patients with nonfluent aphasia, Rezaii et al. (2023) showed that patients with this condition compensate for their deficit in producing complex grammatical structures by increasing the information content of their utterances, dropping high-frequency (function) words in favor of low-frequency (content) words (i.e. “maximizing information”). By modeling variations in grammatical structure based on corpora from 80 languages, Hahn & Xu (2022) found that the resulting models were less optimized for dependency length (DL) and information locality (IL) than the original structures of each language, suggesting that languages evolve to efficiently balance the two considerations of DL and IL. [^20] Notably, Hahn & Xu propose a theory that the interaction between Dependency Length (DL) and Information Locality (IL) accounts for word order variation and change, but neither of the two concepts, as expressed, make claims about directionality/order, only about the relationships between (internal) constituents. In contrast, the Min-Max theory makes the claim that speakers balance information content (as reflected by longer nouns in SV languages and longer verbs in VS languages) with processing/production concerns, and that this interaction is what drives the evolution of word order. As nouns lengthen or verbs shorten in a language, this will affect their placement in a stream, which then impacts the overall linear order observed in surface structures. That is, the Min-Max theory makes a universal claim regarding language-external pressures, while DL and IL are language-internal concepts. These and other studies indicate that language users are sensitive to the pressures of minimizing processing and maximizing information content.
@@
The sensitivity of language users to processing efficiency and information content has also been a major topic of investigation in information theory, which is concerned in part with the encoding and decoding of information streams. One information-theoretic proposal is that semantic systems evolve following an “Information Bottleneck” principle (IB; Tishby et al. 1999), whereby “optimal representations satisfy a tradeoff between compressing X, i.e. minimizing the representational complexity, and maximizing the relevant information about Y” (Zaslavsky 2020: 14). Zaslavsky shows that IB can account for the development of color naming systems in language, and additional research shows the utility of IB in machine learning and distributional models of semantics (Slonim & Tishby 2000).
@@
Implications of this interaction have most recently been examined in agent-based emergent learning simulations with the addition of a “utility” principle (Tucker et al. 2025) – the interaction between “utility”, “informativeness”, and “complexity” allows more human-like communication to emerge among agent models in their study. Such a finding is very much supported by the results reported here, though I would argue that the Min-Max theory offers a more parsimonious account of human behavior as related to language specifically. While the current corpus evidence applies at the word level in relation to basic word order, it seems likely that the minimization of processing difficulty would interact with the maximization of information content at multiple (hierarchical) levels. Speakers min-max in order to select words in a stream (“informativeness” and “complexity” as per Tucker et al.), while simultaneously min-maxing to select relevant structures related to their communicative goals (Tucker et al.’s “utility”).
@@
As noted above, the Min-Max theory also bears some resemblance to Levshina’s (2023) proposals, described in terms of communicative efficiency and accessibility. However, such indicators cannot fully explain certain patterns in grammar such as “harmonic” word order. A recent proposed explanation for such patterns involves frequency of word classes, whereby Mansfield & Krapp (2025) suggest that “harmonic ordering is really a frequency effect: as replication converges on a phrase structure with a consistent linear order, the most frequent word class tends to be at one edge.” Although they did not test token length in their model, it is notable that in Study 1 (§2.1 above) frequency is a major factor in identifying the relationship between length and word order, suggesting that the most frequent word class in their model may also be the longest element of the domain they are testing, which would be driven by the maximization of information principle. The Min-Max theory might also provide an explanation for non-parallel (hierarchical) harmonic word order, which is not explained by their model.
@@
The Min-Max theory may additionally explain why word forms remain long (or lengthen), as described for particular languages (see Cristofaro 2021; Haspelmath 2021; Petré 2017). Here, again, the maximization of information principle would support the development or retention of particular (long) word forms, as speakers balance this need against the minimization of processing difficulty. As noted above, this proposal is supported by evidence from color-naming systems, which show patterns in keeping with IB (Zaslavsky 2020). The Min-Max theory would suggest that this principle does not only hold for “forms” (of, i.e. color words), but also governs “behavior”: speakers behave in ways that are consistent with maximizing information in streams (via word order), while minimizing processing effort.
@@
Such a theory can be conceived of as a universal (underlying) pressure that applies generally to how language is realized. From an innateness perspective, speaker choices impacted by such pressures would lead to a re-analysis or re-formation of rules that guide surface realizations. From a usage-based perspective such pressures would lead to the development of new surface variants. These are not necessarily opposing views (see Rastelli 2025), and the outcomes are quite similarly dependent on language being viewed as a complex adaptive system, whereby diachronic changes may have a combination of both “result-oriented” and “source-oriented” explanations (see Cristofaro 2019; Schmidtke-Bode & Grossman 2019).